# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v2

@@ 4.3 Combinatorial Diagnosis of Different NDs
As shown in Table 6 and 1, there are 19 studies that simultaneously focus on diagnosis of PD, ALS, and HD based on human gait, and 11, 6, 2 studies use C-ML, C-DL, A-DL models, respectively. We provide a brief content summary and discussion for representative studies below.
@@
[101] (Table 6, ID = C-1) uses force sensitive sensors to collect stride intervals and footfall contact times. Features are later derived from these time series data and filtered using a genetic algorithm for diagnosing purpose. The best accuracy results for HC $vs.$ HD, HC $vs.$ PD, HC $vs.$ ALS are achieved using the features selected and SVM models with the corresponding rate of 90.3%, 89.3%, and 96.8% respectively while the accuracy for the diagnosis of three NDDs altogether is 90.6%. [206] (ID = C-2) uses a cross-correlation method to extract features with recurrent neural networks to classify HC and people with neurodegenerative diseases (PD, HD, ALS), achieving highly-competitive performance with 90.6% accuracy. The study of [102] (ID = C-3) exploits 9 statistical features (mean, minimum, etc) of gait rhythm signals ($i.e.$, gait cycle parameters) to perform neurodegenerative disease (ND) diagnosis using different classic machine learning models (SVM, RF, MLP, KNN), and achieves 96.6-100% accuracy. [207] (ID = C-4) proposes a novel analytical methodology for distinguishing healthy individuals and NDs patients based on gait patterns. The method employs intrinsic mode functions (IMFs), and extracts discriminative features from the area measure and relative change in centroid position of the polygon formed by the Convex Hull of these analytic IMFs. Najafabadian $etal.$ [208] (ID = C-5) present a nonlinear gait signal analysis approach to classify NDDs by extracting chaotic and fractal features from gait data and reach a classification accuracy of 92.34%. [209] (ID = C-6) introduces a technique called Statistical Energy Values and Peak Analysis (SEP) for detecting NDDs from signals obtained through force-sensitive resistors. The accuracy, sensitivity, and specificity values are best achieved using the Random Forest Classifier reported at 96.9%, 96.8%, and 96.7% respectively. In [210] (ID = C-7), the authors transform left foot stride and stride interval time series data into a string using symbolic aggregate approximation (SAX). The converted string is then utilized to find a string prototype for each disease using the proposed algorithm called sgUPFCMed. Finally, they implements a fuzzy KNN model to find the best match for a test data sample to classify different diseases and HCs. [211] (ID = C-8) introduces a framework called Topological Motion Analysis (TMA) which embeds gait fluctuation time series into a phase space. In detail, persistent homology is used to extract topological signatures of barcodes, and a Random Forest classifier is employed for analysis. This study demonstrates the effectiveness of TMA in classifying diseases such as ALS, HD, and PD compared to HC, achieving high accuracy of 79.31%, 91.67%, and 87.10% respectively. [212] (ID = C-9) utilizes two feature extraction approaches, $i.e.$, Recurrence Quantification Analysis (RQA) and Fast Walsh-Hadamard Transform (FWHT), to extract gait features and then feed them into different machine learning models such as RF, SVM, and KNN to classify NDs and HC. [213] (ID = C-10) leverages differential transformation of gait force (GF) signals and multiscale sample entropy (MSE) to classify different neurodegenerative diseases using KNN and SVM. This study also shows that the accuracy can be improved by applying the method SMOTE to balance the amount of data in each class. [214] (ID = C-11) proposes a CNN model to differentiate gait of NDs patients and HC based on vGRF signal data (10-second signal) and achieves 95.95-100% accuracy. [103] (ID = C-12) introduces a non-invasive, cost-effective diagnostic tool for analysing gait signals obtained from vGRF sensors. The best classification results are obtained using SVM model and the achieved average accuracy rates are 100% for ALS, 99.78% for PD, and 99.90% for HD. [215] (ID = C-13) explores the usefulness of state space vector, which is combined with KNN, to learn features from gait data for NDD detection. Besides, it also involves statistical features’ extraction from gait signals. It also shows the state space approach’s potential for analyzing other physiological signals. [216] (ID = C-14) builds a preprocessing process, a feature transformation process (PCA), and a classification process (CNN and SVM) to leverage five-minute time-frequency spectrogram of vGRF to classify neurodegenerative diseases. A hybrid model that combines multiple sensor data is proposed in [77] (ID = C-15) to learn gait differences between ALS, HD, PD, and healthy individuals. The model includes a spatial feature extractor (SFE) and a correlative memory neural network (CorrMNN) for capturing representative features and temporal information, which outperforms many state-of-the-art techniques with 99.7-100.0% accuracy. [217] (ID = C-16) proposes to use wavelet coherence to extract features from gait force signal and to represent those extracted features by using spectrogram image. Those features are then processed by CNN to classify HCs and NDs. [218] (ID = C-17) devise Petri net models of human gait with Rs and classification trees (CTs) to classify different neurodegenerative diseases (PD, HD, ALS) from Petri net features of vGRFs, and achieves 91.7-95.1% accuracy. Spatio-temporal gait data are used in [219] (ID = C-18) to measure the energy cost and power spectral density in gait pathologies of different NDs, while fuzzy c-means techniques, learning algorithm for multivariate data analysis (LAMDA), and neural networks are exploited to classify between the neurodegenerative diseases and the control group. In [220] (ID = C-19), the authors compare various deep learning models including LSTM, GRU, InceptionTime, ResNet, FCN, TST, and PatchTST in their ability to differentiate healthy gaits from PD, ALS, and HD gaits.
@@ 4.4 Huntington’s Disease (HD)
As shown in Table 7, there are 7 studies that include AI-assisted HD diagnosis based on human gait (not including 19 combinatorial studies in Table 6). The majority of these studies (6 studies) leverage C-ML models, while other studies (1 study) use C-DL models.
[48] (Table 7, ID = H-1) proposes a method to classify gait between Huntinton’s disease (HD) and health elderly based on IMU data using HMM derived features and SVM classifier, and achieves overall accuracy 90.5% by combining HMM-based features, time and frequency domain features. In [226] (ID = H-2), 11 gait features are extracted from accelerometer data to classify HD patients using different machine learning models (RF, DT, Logiboost, Multibooost) and achieves up to 92.86% accuracy. Two attribute selection algorithms are proposed to choose the most representative gait patterns of subjects. [227] uses pressure data collected during formation of individual footsteps to classify HD by 3D CNN. The experiments showed that using the basic deep learning backbone VGG16 and similar modules can achieve classification accuracy of 89%. Different machine learning models (KNN, SVM, NB, DT, LDA) are utilized in [228] to identify HD from foot force based gait dynamics parameters and achieves up to 100% accuracy with SVM and DT when considering the right foot stance interval. [229] leverages classic machine learning models (SVM, DT, KNN, NB, LDA) to identify HD from the extracted gait features such as right stance interval. In [230], five machine learning models (SVM, DT, NV, RF, LR) are utilized to recognize HD from gait dynamics features and achieves 85.71-97.14% accuracy. [231] uses the IMU data and SVM model to classify HD gait and healthy gait of elderly, achieving 90.5% accuracy.
@@ 4.5 Multiple Sclerosis (MS)
As reported in Table 7, 6 studies that focus on MS diagnosis based on human gait, and there are 4, 1, and 1 using C-ML, C-DL, and A-DL models respectively.
In [105] (Table 7, ID = M-1), the authors uses Extreme Learning Machine (ELM), SVM, MLP, and KNN to classify multiple gait disorders including MS, stroke, and cerebral palsy from gait force data, achieving a accuracy result of 89.8%. [59] (ID = M-2) proposes two normalization methods to normalized the subject’s derived gait characteristics, which are fed into ML models such as SVM, RF, Adaboost, and MLP to classified MS and HC. [232] (ID = M-3) investigates the most efficient gait parameters that could aid in the detection of subtle gait changes in individuals with Multiple Sclerosis. Aiming at this goal, it compares three diagnostic gait systems and employs six different machine learning algorithms. It contributes to the field by providing insights into the use of technology and machine learning in the management of MS. The study of [233] (ID = M-4) devises a CNN-based transfer learning framework to identify MS patients from wearable smartphone sensor data and achieves up to 91.1% accuracy. The interpretations of results also suggest that cadence-based measures, gait speed, and ambulation-related signal perturbations are distinct features for classification. In [106] (ID = M-5), the authors leverage smart-phone and smart-watch to collect gait data and compute related features. Then, the features are feed as input to the ML models (SVM, LR, RF) to classify HC and MS. [58] (ID = M-6) leverages gait parameters extracted from raw data of a walkway system to identify HD patients using different classic machine learning models (SVM, LR, XGBoost) and achieves up to 81% accuracy. The performance is shown to be further improved to 88% when augmenting standard parameters with other custom parameters and normalized subject characteristics.
@@ 4.6 Alzheimer’s Disease (AD)
As presented in Table 7, there are 5 studies that focus on AD diagnosis based on human gait, in which 3, 1, 1 studies exploit C-ML, C-DL, and A-DL models.
In [221] (ID = A-1), the authors explore the use of machine learning classifiers including SVM, MLP, RBNs, and DBSs for diagnosing AD based on postural control kinematics. Four classifiers are compared, and the accuracy of AD diagnosis ranges from 91% to 96.6% when combining postural kinematics and Montreal Cognitive Assessment (MoCA) variable. This study suggests that machine learning models can aid in computer-aided diagnosis of AD using postural control kinematics. [71] (ID = A-2) proposes a cascade neural network approach that combines gait and EEG data for more accurate and efficient classification of AD. The model utilizes attention-based spatial temporal graph convolutional networks (ST-GCN) to extract discriminative features from gait data. Experimental results show superior performance compared to other AD diagnosis methods, with the lower body and right upper limb identified as important for early AD diagnosis. The study of [222] (ID = A-3) collects walking data and derives gait features from participants by using both the ”single-tasking” and ”dual-tasking’ test. The features are then fed as input to the SVM model to classify HC and AD, achieving 78.0% accuracy. In [107] (ID = A-4), the authors explore the feasibility of using the Timed Up and Go (TUG) test, a simple balance and walking assessment, as a tool for detecting AD compared to HC. Joint position data was collected from subjects performing the TUG test using a Kinect camera. Through signal processing and statistical analyses, significant features were identified and used with a support vector machine classifier, achieving high accuracy (97.75%). Different ML models (SVM, KNN, RF, MLP, NB, MLP) [223] (ID = A-5) are assembled to diagnose AD using gait parameters extracted from IMU sensors signals, and achieves about 72.2% accuracy. This study totally proposes seven walking experiment paradigms and collects data using seven wearable devices.
@@ 4.7 Amyotrophic Lateral Sclerosis (ALS)
As reported in Table 7, there are 3 studies that focus on AI-assisted ALS diagnosis based on human gait (excluding 19 combinatorial studies in Table 6), and all studies employ C-ML models.

### v3

@@ 4.3 Combinatorial Diagnosis of Different NDs
As shown in Table 6 and 1, there are 21 studies that simultaneously focus on diagnosis of PD, ALS, and HD based on human gait, and 11, 7, 3 studies use C-ML, C-DL, A-DL models, respectively. We provide a brief content summary and discussion for representative studies below.
@@
Daliri et al. [112] (Table 6, ID = C-1) used force sensitive sensors to collect stride intervals and footfall contact times. Features were later derived from these time series data and filtered using a genetic algorithm for diagnosing purpose. The best accuracy results for HC $vs.$ HD, HC $vs.$ PD, HC $vs.$ ALS were achieved using the features selected and SVM models with the corresponding rate of 90.3%, 89.3%, and 96.8%, respectively. The accuracy for the diagnosis of three NDs altogether was 90.6%. Dutta et al. [215] (ID = C-2) used a cross-correlation method to extract features with recurrent neural networks to classify HC and people with NDs, achieving highly competitive performance with a 90.6% accuracy. The study by Xia et al. [113] (ID = C-3) exploited 9 statistical features (mean, minimum, etc) of gait rhythm signals ($i.e.$, gait cycle parameters) to perform NDs diagnosis using different classic machine learning models (SVM, RF, MLP, KNN), and achieved a 96.6-100% accuracy. Pratiher et al. [216] (ID = C-4) proposed a novel analytical methodology for distinguishing healthy individuals and NDs patients based on gait patterns. The method employed intrinsic mode functions (IMFs), and extracted discriminative features from the area measure and relative change in centroid position of the polygon formed by the Convex Hull of these analytic IMFs. Najafabadian et al. [217] (ID = C-5) presented a nonlinear gait signal analysis approach to classify NDs by extracting chaotic and fractal features from gait data and reached a classification accuracy of 92.34%. Athisakthi [218] (ID = C-6) introduced a technique called Statistical Energy Values and Peak Analysis (SEP) for detecting NDs from signals obtained through force-sensitive resistors. The accuracy, sensitivity, and specificity values were best achieved using the Random Forest Classifier with reported accuracy at 96.9%, 96.8%, and 96.7%, respectively. In the work by Klomsae et al. [219] (ID = C-7), the authors transformed left foot stride and stride interval time series data into a string using symbolic aggregate approximation (SAX). The converted string was then utilized to find a string prototype for each disease using the proposed algorithm called sgUPFCMed. Finally, they implemented a fuzzy KNN model to find the best match for a test data sample to classify different diseases and HCs. Yan et al. [220] (ID = C-8) introduced a framework called Topological Motion Analysis (TMA) which embeds gait fluctuation time series into a phase space. In detail, persistent homology was used to extract topological signatures of barcodes, and a Random Forest classifier was employed for analysis. This study demonstrated the effectiveness of TMA in classifying diseases such as ALS, HD, and PD compared to HC, achieving high accuracies of 79.31%, 91.67%, and 87.10%, respectively. Begum et al. [221] (ID = C-9) utilized two feature extraction approaches, $i.e.$, Recurrence Quantification Analysis and Fast Walsh-Hadamard Transform, to extract gait features and then fed them into different ML models such as RF, SVM, and KNN to classify NDs and HC. Nam et al. [222] (ID = C-10) leveraged differential transformation of gait force (GF) signals and multiscale sample entropy (MSE) to classify different neurodegenerative diseases using KNN and SVM. This study also showed that the accuracy could be improved by applying the method SMOTE to balance the amount of data in each class. Lin et al. [223] (ID = C-11) proposed a CNN model to differentiate gait of NDs patients and HC based on vGRF signal data (10-second signal) and achieved a 95.95-100% accuracy. Beyrami et al. [114] (ID = C-12) introduced a non-invasive, cost-effective diagnostic tool for analysing gait signals obtained from vGRF sensors. The best classification results were obtained using SVM model and the achieved average accuracy rates were 100% for ALS, 99.78% for PD, and 99.90% for HD. Yilmaz et al. [224] (ID = C-13) explored the usefulness of state space vector, which is combined with KNN, to learn features from gait data for NDs detection. Besides, it also involved statistical features’ extraction from gait signals. It also showed the state space approach’s potential for analyzing other physiological signals. Setiawan et al. [225] (ID = C-14) builded a preprocessing process, a feature transformation process (PCA), and a classification process (CNN and SVM) to leverage five-minute time-frequency spectrogram of vGRF to classify NDs. A hybrid model that combined multiple sensor data was proposed in Zhao et al. [82] (ID = C-15) to learn gait differences between ALS, HD, PD, and healthy individuals. The model included a spatial feature extractor (SFE) and a correlative memory neural network (CorrMNN) for capturing representative features and temporal information, which outperformed many state-of-the-art techniques with a 99.7-100.0% accuracy. Setiawan et al. [226] (ID = C-16) proposed to use wavelet coherence to extract features from gait force signal and to represent those extracted features by using spectrogram image. Those features were then processed by CNN to classify HCs and NDs. Tobar et al. [227] (ID = C-17) devised Petri net models of human gait with Rs and classification trees (CTs) to classify different NDs(e.g., PD, HD, ALS) from Petri net features of vGRFs, and achieved a 91.7-95.1% accuracy. Spatio-temporal gait data were used in the work by Torres et al. [228] (ID = C-18) to measure the energy cost and power spectral density in gait pathologies of different NDs, while fuzzy c-means techniques, learning algorithm for multivariate data analysis (LAMDA), and neural networks are exploited to classify between NDs and the control group. The authors compared various deep learning models including LSTM, GRU, InceptionTime, ResNet, FCN, TST, and PatchTST in their ability to differentiate healthy gaits from PD, ALS, and HD gaits (Zhou et al. [229] (ID = C-19)).
@@ 4.4 Huntington’s Disease (HD)
As shown in Table 7, there are 7 studies that include AI-assisted HD diagnosis based on human gait (not including 21 combinatorial studies in Table 6). The majority of these studies (6 studies) leverage C-ML models, while other studies (1 study) use C-DL models.
Mannini et al. [53] (Table 7, ID = H-1) proposed a method to classify gait between HD and health elderly based on IMU data using HMM derived features and SVM classifier, and achieved an overall accuracy of 90.5% by combining HMM-based , time, and frequency domain features. Eleven gait features were extracted from accelerometer data to classify HD patients using different machine learning models (e.g., RF, DT, Logiboost, Multibooost) achieving up to 92.86% accuracy. Two attribute selection algorithms were proposed to choose the most representative gait patterns of subjects (Acosta et al. [238] (ID = H-2)). Zhang et al. [239] (ID = H-3) used pressure data collected during formation of individual footsteps to classify HD by 3D CNN. The experiments showed that using the basic deep learning backbone VGG16 and similar modules could achieve a classification accuracy of 89%. Different machine learning models (KNN, SVM, NB, DT, LDA) were utilized in the work by Felix et al. [240] (ID = H-4) to identify HD from foot force based gait dynamics parameters and achieved up to 100% accuracy with SVM and DT when considering the right foot stance interval. Paula et al. [241] (ID = H-5) leveraged classic machine learning models (SVM, DT, KNN, NB, LDA) to identify HD from the extracted gait features such as right stance interval. Huang et al. [242] (ID = H-6) utilized five machine learning models (SVM, DT, NV, RF, LR) to recognize HD from gait dynamics features and achieved 85.71-97.14% accuracy. Trojaniello [243] (ID = H-7) used the IMU data and SVM model to classify HD gait and healthy gait of elderly, achieving a 90.5% accuracy.
@@ 4.5 Multiple Sclerosis (MS)
As reported in Table 7, 6 studies that focus on MS diagnosis based on human gait, and there are 4, 1, and 1 using C-ML, C-DL, and A-DL models, respectively.
Patil et al. [116] (Table 7, ID = M-1), used Extreme Learning Machine (ELM), SVM, MLP, and KNN to classify multiple gait disorders including MS, stroke, and cerebral palsy from gait force data, achieving an accuracy result of 89.8%. Kaur et al. [64] (ID = M-2) proposed two normalization methods to normalized the subject’s derived gait characteristics, which were fed into ML models such as SVM, RF, Adaboost, and MLP to classify MS and HC. Trentzsch et al. [244] (ID = M-3) investigated the most efficient gait parameters that could aid in the detection of subtle gait changes in individuals with MS. Aiming at this goal, they compared three diagnostic gait systems and employed six different machine learning algorithms. They contributed to the field by providing insights into the use of technology and machine learning in the management of MS. The study of Creagh et al. [245] (ID = M-4) devised a CNN-based transfer learning framework to identify MS patients from wearable smartphone sensor data and achieved up to 91.1% accuracy. The interpretations of results also suggested that cadence-based measures, gait speed, and ambulation-related signal perturbations were distinct features for classification. The work by Creagh et al. [117] (ID = M-5) leveraged smart-phone and smart-watch to collect gait data and compute related features. Then, the features were fed as input to the ML models (e.g., SVM, LR, RF) to classify HC and MS. Hu et al. [63] (ID = M-6) leveraged gait parameters extracted from raw data of a walkway system to identify HD patients using different classic ML models (e.g., SVM, LR, XGBoost) and achieved up to 81% accuracy. The performance was shown to be further improved to 88% when augmenting standard parameters with other custom parameters and normalized subject characteristics.
@@ 4.6 Alzheimer’s Disease (AD)
As presented in Table 7, there are 6 studies that focus on AD diagnosis based on human gait, in which 3, 1, 2 studies exploit C-ML, C-DL, and A-DL models.
In the work by Costa et al. [232] (ID = A-1), the authors explored the use of ML classifiers including SVM, MLP, RBNs, and DBSs for diagnosing AD based on postural control kinematics. Four classifiers were compared, and the accuracy of AD diagnosis ranges from 91% to 96.6% when combining postural kinematics and Montreal Cognitive Assessment (MoCA) variable. This study suggested that machine learning models could aid in computer-aided diagnosis of AD using postural control kinematics. You et al. [76] (ID = A-2) proposed a cascade neural network approach that combined gait and EEG data for more accurate and efficient classification of AD. The model utilized attention-based ST-GCN to extract discriminative features from gait data. Experimental results showed superior performance compared to other AD diagnosis methods, with the lower body and right upper limb identified as important for early AD diagnosis. The study of Ghoraani et al. [233] (ID = A-3) collected walking data and derived gait features from participants by using both the ”single-tasking” and ”dual-tasking’ test. The features were then fed as input to the SVM model to classify HC and AD, achieving a 78.0% accuracy. In the work by Seifallahi et al. [118] (ID = A-4), the authors explored the feasibility of using the Timed Up and Go (TUG) test, a simple balance and walking assessment, as a tool for detecting AD from HC. Joint position data was collected from subjects performing the TUG test using a Kinect camera. Through signal processing and statistical analyses, significant features were identified and used with a support vector machine classifier, achieving a high accuracy (97.75%). Different ML models (SVM, KNN, RF, MLP, NB, MLP) by Jeon et al. [234] (ID = A-5) were assembled to diagnose AD using gait parameters extracted from IMU sensors signals, and achieved about 72.2% accuracy. This study totally proposed seven walking experiment paradigms and collected data using seven wearable devices.
@@ 4.7 Amyotrophic Lateral Sclerosis (ALS)
As reported in Table 7, there are 3 studies that focus on AI-assisted ALS diagnosis based on human gait (excluding 21 combinatorial studies in Table 6), and all studies employ C-ML models.